An engineer designed a valve that will regulate water pressure on an automobile engine. The engineer designed the valve such that it would produce a mean pressure of 5.4 pounds/square inch. It is believed that the valve performs above the specifications. The valve was tested on 9 engines and the mean pressure was 5.7 pounds/square inch with a variance of 0.81. A level of significance of 0.025 will be used. Assume the population distribution is approximately normal. Make the decision to reject or fail to reject the null hypothesis.
Answer #1

n=9, X= 5.7 ² = 0.81 Ho Ho: l=5.4 el s. 4 t = X- el = 5.7-5.4 0.81 rg t = 1 2=0.025 : CR = t 8,0.025 = 2.306 It LCR. we fail